OpenStudy (anonymous):

Which of the following is the slope between the points (-2, 4) and (2, -2)? negative three halves negative two thirds three halves two thirds

OpenStudy (tkhunny):

Have you considered simply calculating that slope?

OpenStudy (anonymous):

good idea

OpenStudy (tkhunny):

Great! Let's see it.

OpenStudy (anonymous):

i got 0,-2? thats wrong and i know it

OpenStudy (anonymous):

help please

OpenStudy (tkhunny):

\(Slope = \dfrac{Change\;in\;y}{Change\;in\;x}=\dfrac{y_{2} - y_{1}}{x_{2}-x_{1}} = \dfrac{-2-4}{2-(-2)}\)

OpenStudy (anonymous):

-3,2?

OpenStudy (tkhunny):

-6/4 = -3/2 Not certain about your use of the comma ",". Does that mean division?

OpenStudy (anonymous):

yes so i got it right thx! :) one more question please?

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, -2) and perpendicular to y = one fourthx + 2.

OpenStudy (tkhunny):

Where does a comma mean division. I've never seen that usage? What book? What country? What culture?

OpenStudy (anonymous):

which one y = 4x - 6 y = -4x - 6 y = -4x + 2 y = 4x

OpenStudy (tkhunny):

Do you have the Point-Slope form?

OpenStudy (anonymous):

look above for question

OpenStudy (tkhunny):

I saw the problem statement. Now, answer my question.

OpenStudy (anonymous):

dont quite understand jus help

OpenStudy (tkhunny):

I am helping if you answer my question. There are various "forms" for the equation of a line... Point-Slope Slope-Intercept 2-Point Intercept Which of these are familiar to you?

OpenStudy (anonymous):

slope interept

OpenStudy (tkhunny):

Awesome. Well, you're going to need the Point-Slope for this one. It is this \((y - y_{0}) = m(x - x_{0})\) m is the same as y = mx + b. It is just the slope of the line. The point \((x_{0},y_{0})\) is just any point on the line. This is just barely different from the slope-intercept form. With slope-intercept, we know the point \((0,b)\) - the y-intercept. With point-slope, it can be ANY point. So, where does that leave us?

OpenStudy (anonymous):

the 3rd one?

OpenStudy (tkhunny):

Choose the equation of the line passing through the point (-1, -2) and perpendicular to y = one fourthx + 2. I presume this means: Containing the point \((-1,-2)\) Slope perpendicular to \(y = ¼x + 2\) Okay, how did you get your answer?

OpenStudy (anonymous):

look i dont have all day can you just give me a hint to the answer

OpenStudy (tkhunny):

I give you hint after hint and you ask for a hint? I'm here for you to learn. I'm a volunteer. If you don't want to learn, you are wasting my time. By definition, I cannot possibly be wasting your time. you have a point-slope form \((y-y_{0})=m(x-x_{0})\) You have a point that belongs to your line. \((-1,-2)\) This gives \((y+2)=m(x+1)\) What is the slope of your new line?